# A shared-revenue Bertrand game

### v1

@@ 6.2. Refined equilibria of outside option game
– an admissible, Pareto optimal equilibrium strategy profile of the shared-revenue Bertrand game for every set of parameters
@@
If the retailer chooses a fee $\alpha>\alpha_{\max}$, they will achieve a payoff of 00 in equilibrium of the Outside Option Game because the independent seller will leave. The independent seller, on the other hand, will achieve a payoff of $\pi_{s}^{(\ell)}(\delta)$ for $\alpha>\alpha_{\max}$, which is therefore a lower bound on their equilibrium payoff. Additionally, in equilibrium of the Outside Option Game the retailer will choose a fee $\alpha\leq\alpha_{\max}$, which bounds the maximum fee further away from $1$.
@@
The right subplot of Figure 5 shows another interesting case where $\delta$ is low enough such that the independent seller achieves a higher payoff than the retailer in equilibrium. In fact, the independent seller is so competitive with the retailer that the retailer achieves a significantly lower payoff in equilibrium of the Outside Option Game than they would if the indepdendent seller did not exist. Additionally, notice that $\alpha_{\max}<\alpha^{\dagger}_{r}$ is not in the continuum region. Because of this, the choice of $\rho$ does not matter, which should be clear graphically because the entire continuum region is “grayed out” for both the retailer and independent seller. In fact, we can say more generally that
(34) $\alpha^{(o)}_{*}=\min\left\{\alpha_{*},\alpha_{\max}\right\}\quad\text{ if }\alpha_{\max}\leq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$
@@
The Outside Option Game is the final extension of the shared-revenue Bertrand game that we explore in this paper. The game captures the effect that the cost configurations ($c_{r},c_{s}$) and the credibility of the independent seller’s outside option threat ($\delta$) have on the optimal fee structures and the minimum/maximum payoffs achievable in equilibrium of the revenue sharing program.
@@ 7. Conclusion and future work
In this work, we presented a game theory model of a shared-revenue Bertrand game to analyze the dynamics and incentives induced by revenue sharing programs. Furthermore, we investigated some natural extensions of the model to make it more realistic by modeling the proportion of revenue sharing as an action and giving the independent seller an outside option to sell on their own. Our findings resonate strongly with existing results from duopolistic competition and optimal taxation theory. In equilibrium, both players are able to achieve positive payoff, and the referral fee is set at a “sweet spot” low enough to stimulate economic activity while high enough to maximize payoff.
@@
  • To improve the present analysis of the shared-revenue Bertrand game, some natural next steps would be to more completely consider general demand curves and mixed strategies. While we anticipate qualitatively similar equilibria under appropriate assumptions on the demand curves and the mixing distributions, there is much left to formally prove this.
@@
  • Empirical studies of revenue sharing programs could reveal aspects of the real world our models fail to capture, including in the other application domains introduced in Section 1.1.
@@ Appendix A Notation table
$q_{x}(p_{r},p_{s})$ (5) Quantity demanded for player $x$
@@
$\rho(\alpha,c_{r},c_{s})$ $\in\mathcal{S}_{p}$ Equilibrium strategy profile of shared-revenue Bertrand game
@@
$\pi_{x}^{\text{(eq)}}(\rho,\alpha,c_{r},c_{s})$ $\in[0,1)$ Payoff in equilibrium of shared-revenue Bertrand game
@@
$c_{s^{*}}$ $\frac{2c_{r}^{2}}{1+c_{r}}$ Seller optimal feasibility cost; $c_{s}\leq c_{s^{*}}$ is necessary for $(p_{s^{*}},s)$ to be a potential equilibrium
$p_{s^{*}}$ $\frac{1}{2}\left(1+\frac{c_{s}}{1-\alpha}\right)$ Seller optimal single agent price, the maximizer of $\pi_{s,s}$
$p_{s^{*}}^{(\ell)}$ $\frac{1}{2}\left(1+c_{s}+\delta\right)$ Seller optimal single agent price in leaving subgame
$p_{r^{*}}$ $\frac{1}{2}\left(1+c_{r}\right)$ Retailer optimal single agent price, the maximizer of $\pi_{r,r}$
@@
$p^{\dagger}$ (12) Retailer optimality equivalence price, satisfying $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$
$\alpha^{\dagger}_{s}$ (62) Seller $p^{\dagger}$ relevance fee, below which selling at $p^{\dagger}$ becomes a potential best response
$\alpha^{\dagger}_{r}$ $\frac{1-c_{r}}{1+c_{r}}$ Retailer $p^{\dagger}$ relevance fee, above which allowing the independent seller to sell at $p^{\dagger}$ becomes a potential best response
$\alpha_{s^{*}}$ $1-2c_{r}+c_{s}$ Seller optimal price feasibility fee, above which $(p_{s^{*}},s)$ becomes a potential equilibrium
$\alpha_{r^{*}}$ $1-\frac{2c_{s}}{1+c_{r}}$ Retailer optimal price feasibility fee, above which $(p_{r^{*}},r)$ becomes a potential equilibrium
$\alpha_{s^{*},r^{*}}$ $\frac{c_{r}-c_{s}}{c_{r}}$ Optimality switching fee satisfying $p_{s^{*}}(\alpha_{s^{*},r^{*}})=p_{r^{*}}$
@@
$\overline{\alpha}$ See Lemma 4 Retailer equilibrium fee with strategy profile $\overline{\rho}$
$\underline{\alpha}$ $\alpha_{r^{*}}$ Retailer equilibrium fee with strategy profile $\underline{\rho}$
$\alpha^{(o)}_{*}$ (32) Retailer equilibrium fee for the Outside Option Game
$\alpha_{\max}$ (33) Maximum fee for which the independent seller prefers to stay in the revenue sharing program rather than exercise their outside option in the Outside Option Game
@@ B.1. The retailer and seller can achieve nonzero payoff
In a shared-revenue Bertrand game with $c_{r}\geq 1$, the only admissible Nash equilibrium is $(p_{r}=1,p_{s}=\min\left\{p_{s^{*}},1\right\})$.
@@ B.2. Independent seller’s cost is lower
In a shared-revenue Bertrand game with $c_{r}<c_{s}$, there are no Nash equilibria where the independent seller fulfills demand.
@@
Appendix C Proof of Proposition 1
@@
Let $(p_{r},p_{s})=(p,p)$, with the equilibrium quantities sold by each player $q_{r}=\beta(1-p)$ and $q_{s}=(1-\beta)(1-p)$. Then $\beta\in\{0,1\}$.
@@
Let $p$ be an equilibrium market price, with the equilibrium quantities sold by each player $q_{r}=\beta(1-p)$ and $q_{s}=(1-\beta)(1-p)$, with $\beta\notin\{0,1\}$. Then the total quantity sold is $q=q_{r}+q_{s}=1-p$ and the retailer’s payoff is
(35) $\pi_{r}$ $=(p-c_{r})q_{r}+\alpha pq_{s}$
$=(p-c_{r})\beta(1-p)+\alpha p(1-\beta)(1-p)$
$=\left[(p-c_{r})\beta+\alpha p(1-\beta)\right](1-p).$
We examine three possible cases — $(p-c_{r})>\alpha p$, $(p-c_{r})<\alpha p$ and $(p-c_{r})=\alpha p$ — and show that in each case, one of the players has an incentive to deviate to a different price.
If $(p-c_{r})>\alpha p$ (the retailer’s payoff per unit it sells is larger than its payoff from the independent seller fullfilling that unit), then the retailer prefers to deviate by setting a price $p-\varepsilon$ to take the entire market. Similarly, if $(p-c_{r})<\alpha p$, the retailer can achieve a higher payoff by deviating to price $p+\varepsilon$, allowing the the independent seller to fulfill the demand. When $(p-c_{r})=\alpha p$, the independent seller’s payoff is $q_{s}=(1-\beta)(p-c_{s}-\alpha p)$. A similar argument, comparing $(p-c_{s}-\alpha p)$ and 00, shows that the independent seller will either decrease its price to take the entire market, or increase its price and allow the retailer to fulfill the demand (since the independent seller only achieves nonzero payoff when they fulfills demand themselves, this last case will only happen if $p$ is below the independent seller’s breakeven price). Note that the case of $(p-c_{r})=\alpha p$ and $(p-c_{s}-\alpha p)=0$ is impossible, as that would imply $c_{r}=c_{s}$, which contradicts our assumption that $c_{r}>c_{s}$. Thus, if $\beta\notin\{0,1\}$, the proposed price and market split does not constitute an equilibrium. ∎

### v2

@@ 6.2. Refined equilibria of shared-revenue Bertand game
– an admissible, Pareto optimal equilibrium strategy profile of the staying subgame for every set of parameters
@@
If the retailer chooses a fee $\alpha>\alpha_{\max}$, they will achieve a payoff of 0 in equilibrium of the shared-revenue Bertrand game because the independent seller will leave. The independent seller, on the other hand, will achieve a payoff of $\pi_{s}^{(\ell)}(\delta)$ for $\alpha>\alpha_{\max}$, which is therefore a lower bound on their equilibrium payoff. Additionally, in equilibrium of the shared-revenue Bertrand game the retailer will choose a fee $\alpha\leq\alpha_{\max}$, which bounds the maximum fee further away from $1$.
@@
The right subplot of Figure 5 shows another interesting case where $\delta$ is low enough such that the independent seller achieves a higher payoff than the retailer in equilibrium. In fact, the independent seller is so competitive with the retailer that the retailer achieves a significantly lower payoff in equilibrium of the shared-revenue Bertrand game than they would if the independent seller did not exist. Additionally, notice that $\alpha_{\max}<\alpha^{\dagger}_{r}$ is not in the continuum region. Because of this, the choice of $\rho$ does not matter, which should be clear graphically because the entire continuum region is “grayed out” for both the retailer and independent seller. In fact, we can say more generally that
(30) $\alpha^{(o)}_{*}=\min\left\{\alpha_{*},\alpha_{\max}\right\}\quad\text{ if }\alpha_{\max}\leq\min\left\{\alpha^{\dagger}_{r},\alpha_{s^{*}}\right\}$
@@ 7. Conclusion and future work
In this work, we presented a game theory model of a shared-revenue Bertrand game to analyze the dynamics and incentives induced by revenue sharing programs. We model the revenue sharing program as a negotiation where the retailer chooses the referral fee then the independent seller chooses whether to participate or leave and sell on their own. Our findings resonate strongly with existing results from duopolistic competition and optimal taxation theory. In equilibrium, both players are able to achieve positive payoff, and the referral fee is set low enough to stimulate economic activity while high enough to maximize payoff.
@@ Appendix A Notation table
$q(p)$ $\in[0,1]$ Quantity demanded at price $p$
@@
$\rho(\alpha,c_{r},c_{s})$ $\in\mathcal{S}_{p}$ Equilibrium strategy profile of staying subgame
@@
$\pi_{x}^{\text{(eq)}}(\rho,\alpha,c_{r},c_{s})$ $\in[0,1)$ Payoff in equilibrium of staying subgame
@@
$c_{s^{*}}$ $\frac{c_{r}^{2}}{p_{r^{*}}}$ Seller optimal feasibility cost; $c_{s}\leq c_{s^{*}}$ is necessary for $(p_{s^{*}},s)$ to be a potential equilibrium
$p_{s^{*}}$ $\arg\max_{p}\pi_{s,s}(p)$ Seller optimal single agent price, the maximizer of $\pi_{s,s}$
$p_{s^{*}}^{(\ell)}$ Seller optimal single agent price in leaving subgame
$p_{r^{*}}$ $\arg\max_{p}\pi_{r,r}(p)$ Retailer optimal single agent price, the maximizer of $\pi_{r,r}$
@@
$p^{\dagger}$ Retailer optimality equivalence price with $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$
$\tilde{p}$ $\arg\max_{p}\pi_{r,s}(p)$ Revenue maximizing price
$\alpha^{\dagger}_{s}$ Seller $p^{\dagger}$ relevance fee, below which selling at $p^{\dagger}$ becomes a potential best response
$\alpha^{\dagger}_{r}$ $1-\frac{c_{r}}{p_{r^{*}}}$ Retailer $p^{\dagger}$ relevance fee, above which allowing the independent seller to sell at $p^{\dagger}$ becomes a potential best response
$\alpha_{s^{*}}$ $1-\frac{c_{r}}{p_{s^{*}}}$ Seller optimal price feasibility fee, above which $(p_{s^{*}},s)$ becomes a potential equilibrium
$\alpha_{r^{*}}$ $1-\frac{c_{s}}{p_{r^{*}}}$ Retailer optimal price feasibility fee, above which $(p_{r^{*}},r)$ becomes a potential equilibrium
$\alpha_{s^{*},r^{*}}$ $1-\frac{c_{s}}{c_{r}}$ Optimality switching fee satisfying $p_{s^{*}}(\alpha_{s^{*},r^{*}})=p_{r^{*}}$
@@
$\overline{\alpha}$ (25) Retailer equilibrium fee with strategy profile $\overline{\rho}$
$\alpha^{(o)}_{*}$ (28) Retailer equilibrium fee for the shared-revenue Bertrand game
$\alpha_{\max}$ (29) Maximum fee for which the independent seller prefers to stay in the revenue sharing program rather than exercise their outside option in the shared-revenue Bertrand game
@@ B.1. The retailer and seller can achieve nonzero payoff
In a staying subgame with $c_{r}\geq 1$, the only admissible Nash equilibrium is $(p_{r}=1,p_{s}=\min\left\{p_{s^{*}},1\right\})$.
@@ B.2. Independent seller’s cost is lower
In a staying subgame with $c_{r}<c_{s}$, there are no Nash equilibria where the independent seller fulfills demand.
@@
Appendix C Key prices and lemmas
C.1. No split market equilibria
In this section, we present the full proof of Proposition 1, reprised here for convenience –
@@ C.1. No split market equilibria
Let $(p_{r},p_{s})=(p,p)$, with the equilibrium quantities sold by each player $q_{r}=\beta q(p)$ and $q_{s}=(1-\beta)q(p)$. Then $\beta\in\{0,1\}$.
@@
Suppose for contradiction $\beta\in(0,1)$. For such an equilibrium, the retailer’s payoff is
$\pi_{r}(p)=\beta\pi_{r,r}(p)+(1-\beta)\pi_{r,s}(p)$
If $p<p_{r_{\text{ind}}}$, then the retailer clearly prefers to increase their price to $p+\varepsilon$ and allow the independent seller to fulfill demand at price $p$, achieving payoff $\pi_{r,s}(p)$ rather than the weighted average.
If $p>p_{r_{\text{ind}}}$, then $\pi_{r,r}(p)>\pi_{r,s}(p)$, so by continuity of $q(p)$ there exists some $\varepsilon>0$ such that the retailer prefers to achieve payoff $\pi_{r}(p-\varepsilon)=\pi_{r,r}(p-\varepsilon)$. Specifically, suppose that $\pi_{r,r}(p-\varepsilon)=\pi_{r,r}(p)-\delta$. Indeed, we need to find a $\delta$ satisfying
$\pi_{r,r}(p)-\delta=\pi_{r,r}(p-\varepsilon)>\beta\pi_{r,r}(p)+(1-\beta)\pi_{r,s}(p)\iff\delta<(1-\beta)(\pi_{r,r}(p)-\pi_{r,s}(p))$
which always exists by continuity of $\pi_{r,r}(p)$ and the fact that $\pi_{r,r}(p)-\pi_{r,s}(p)>0$.
Finally, if $p=p_{r_{\text{ind}}}$, we have $p>p_{s_{\text{ind}}}$, so a similar argument as the previous case for the independent seller implies that they would prefer to increase their price to $p-\varepsilon$, allowing them to achieve a payoff of
$\pi_{r,s}(p-\varepsilon)>(1-\beta)\pi_{r,s}(p)$
by continuity of $\pi_{r,s}(p)$. Thus, if $\beta\notin\{0,1\}$, the proposed price and market split does not constitute an equilibrium. ∎
C.2. Payoff function properties
First, note that the revenue function $pq(p)$ is strictly concave because
$\frac{\partial^{2}}{\partial p^{2}}pq(p)=\frac{\partial}{\partial p}(q(p)+pq^{\prime}(p))=(1+p)q^{\prime\prime}(p)+q^{\prime}(p)<0$
thus, it is uniquely maximized on $(0,1)$ and the price $\tilde{p}\equiv\arg\max_{p}pq(p)$ is well-defined. Notice $\tilde{p}$ is trivially also the maximizer of $\pi_{r,s}(p)=(1-\alpha)pq(p)$. For the payoff function $\pi_{r,r}(p)=pq(p)-c_{r}q(p)$, using the above result we have
$\frac{\partial^{2}}{\partial p^{2}}\pi_{r,r}(p)=\frac{\partial^{2}}{\partial p^{2}}(pq(p)-c_{r}q(p))=(1-c_{r}+p)q^{\prime\prime}(p)+q^{\prime}(p)$
and since $c_{r}<1$, $\pi_{r,r}$ is also strictly concave and uniquely maximized which implies $p_{r^{*}}$ is well-defined. Similarly, $\pi_{s,s}(p)$ is strictly concave for all prices whenever $c_{s}\leq 1-\alpha$, so $p_{s^{*}}$ is well-defined. For $c_{s}>1-\alpha$, there is no price at which the independent seller can achieve positive payoff, so $p_{s^{*}}$ will not be relevant. We now remark about the relative ordering of these maximizers –
Lemma 0.
Let $\tilde{p}\equiv\arg\max_{p}\pi_{r,s}(p)$. Then, $p_{r^{*}},p_{s^{*}},p^{\dagger}>\tilde{p}$, and $p_{s^{*}}>p_{s_{\text{ind}}}$ for $\alpha<1-c_{s}$.
Proof.
Consider the first order conditions satisfied by $\tilde{p}$ and $p_{r^{*}}$
(31) $\frac{q^{\prime}(\tilde{p})}{q(\tilde{p})}$ $=-\frac{1}{\tilde{p}}$
(32) $\frac{q^{\prime}(p_{r^{*}})}{q(p_{r^{*}})}$ $=\frac{1}{c_{r}-p_{r^{*}}}$
Notice that the left side is the same in both equations, while the function on the right side is shifted by $c_{r}$. Thus, the price at which the two functions intersect also shifts right, implying that $p_{r^{*}}>\tilde{p}$. Now, consider the first order condition for $p_{s^{*}}$
(33) $\frac{q^{\prime}(p_{s^{*}})}{q(p_{s^{*}})}=\frac{1}{\frac{c_{s}}{1-\alpha}-p_{s^{*}}}$
Clearly, the same argument suffices to imply $p_{s^{*}}>\tilde{p}$. Furthermore, notice that $\frac{q^{\prime}(p)}{q(p)}<0$, so we must have $p_{s^{*}}>\frac{c_{s}}{1-\alpha}=p_{s_{\text{ind}}}$ in order for the right side to be negative as well and for there to be an intersection. However, when $\alpha\geq 1-c_{s}\iff p_{s_{\text{ind}}}\geq 1$, the independent seller cannot achieve positive payoff at any price, thus their optimal price would be $p_{s^{*}}=1\leq p_{s_{\text{ind}}}$ (because $q(1)=\pi_{s,s}(1)=0$ per our normalization choices).
Finally, we turn our attention to showing $p^{\dagger}>\tilde{p}$. The only nontrivial case is when $p^{\dagger}<\min\{p_{s^{*}},p_{r^{*}}\}$. Recall that $p^{\dagger}$ was defined as the greatest price such that $\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$. Suppose for contadiction $p^{\dagger}<\tilde{p}$; this implies $\pi_{r,s}(\tilde{p})>\pi_{r,r}(p_{r^{*}})$. However, since $\pi_{r,s}$ is continuous with $\pi_{r,s}(1)=0$, this implies there must exist some ${p^{\dagger}}^{\prime}\in(\tilde{p},1)$ such that $\pi_{r,s}({p^{\dagger}}^{\prime})=\pi_{r,r}(p_{r^{*}})$. However, this implies ${p^{\dagger}}^{\prime}>\tilde{p}>p^{\dagger}$, contradicting the maximality of $p^{\dagger}$ and concluding the proof. ∎
Corollary 0.
The payoff functions $\pi_{r,r}$, $\pi_{s,s}$, and $\pi_{r,s}$ are unimodal on $[0,1]$, in the sense that they have only one critical point.
Proof.
This follows from the first order conditions above; in each, the function on the right side is strictly increasing in $p$, and the function on the left is strictly decreasing in $p$ because
$\frac{\partial}{\partial p}\left(\frac{q^{\prime}(p)}{q(p)}\right)=\frac{q(p)q^{\prime\prime}(p)-(q^{\prime}(p))^{2}}{q^{2}(p)}<0$
Thus, the intersection price must be unique which is equivalent to the payoff function being unimodal. ∎
C.3. Key prices and relative ordering
In this section we derive the key prices and fees introduced in Sections 2.2 and 3.3. We begin with the indifference prices –
$\pi_{r,r}(p_{r_{\text{ind}}})$ $=\pi_{r,s}(p_{r_{\text{ind}}})$
$(p_{r_{\text{ind}}}-c_{r})q(p_{r_{\text{ind}}})$ $=\alpha p_{r_{\text{ind}}}q(p_{r_{\text{ind}}})$
(34) $p_{r_{\text{ind}}}$ $=\frac{c_{r}}{1-\alpha}$
$\pi_{s,s}(p_{s_{\text{ind}}})$ $=\pi_{s,r}(p_{s_{\text{ind}}})$
$((1-\alpha)p_{s_{\text{ind}}}-c_{s})q(p_{s_{\text{ind}}})$ $=0$
(35) $p_{s_{\text{ind}}}$ $=\frac{c_{s}}{1-\alpha}$
Next, we will derive the fees that demarcate the changes in ordering between the prices. Note that there is no closed form for $\alpha^{\dagger}_{s}$. The simplest are $\alpha_{r^{*}}$ and $\alpha_{s^{*}}$
(36) $p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}\leq p_{r^{*}}\iff 1-\alpha\geq\frac{c_{s}}{p_{r^{*}}}\iff\alpha\leq 1-\frac{c_{s}}{p_{r^{*}}}\equiv\alpha_{r^{*}}$
(37) $p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}\leq p_{s^{*}}\iff 1-\alpha\geq\frac{c_{s}}{p_{s^{*}}}\iff\alpha\leq 1-\frac{c_{s}}{p_{s^{*}}}\equiv\alpha_{s^{*}}$
Next we find an expression for $\alpha_{s^{*},r^{*}}$. Indeed, considering the functional form of the first order conditions (32) and (33), we clearly have
(38) $p_{r^{*}}=p_{s^{*}}\iff\frac{c_{s}}{1-\alpha_{s^{*},r^{*}}}=c_{r}\iff\alpha_{s^{*},r^{*}}=1-\frac{c_{s}}{c_{r}}$
To determine the ordering, note that as $\alpha\downarrow 0$, $p_{s_{\text{ind}}}\downarrow c_{s}$, which implies by the first order conditions that $p_{s^{*}}\leq p_{r^{*}}$ when $\alpha=0$. Thus, the ordering must be $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$.
Finally, it remains to find a closed form for $\alpha^{\dagger}_{r}$. Note that $\alpha^{\dagger}_{r}$ determines the relative ordering of both $p_{r_{\text{ind}}}$ and $p_{r^{*}}$ as well as $p^{\dagger}$ and $p_{r^{*}}$. Thus in order to show $\alpha^{\dagger}_{r}$ is well-defined, we begin by proving the following lemma –
Lemma 0.
$p_{r_{\text{ind}}}\leq p_{r^{*}}\iff p^{\dagger}\leq p_{r^{*}}$.
Proof.
For the forward direction suppose $p_{r_{\text{ind}}}\leq p_{r^{*}}$ and suppose for contradiction $p^{\dagger}>p_{r^{*}}\geq p_{r_{\text{ind}}}$. This implies $\pi_{r,r}(p^{\dagger})>\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$. However, this is clearly a contradiction since $p_{r^{*}}$ is the unique maximizer of $\pi_{r,r}$.
For the reverse direction, suppose $p^{\dagger}\leq p_{r^{*}}$ and suppose for contradiction $p_{r_{\text{ind}}}>p_{r^{*}}$. Note $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$, and since we have $\tilde{p}\leq p^{\dagger}\leq p_{r^{*}}$ by Lemma 2, it follows that
$\pi_{r,s}(p_{r^{*}})\leq\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$
However, this contradicts the definition of $p_{r_{\text{ind}}}$ since $p_{r^{*}}<p_{r_{\text{ind}}}$ but $\pi_{r,s}(p_{r^{*}})\leq\pi_{r,r}(p_{r^{*}})$. This concludes the proof of the reverse direction. ∎
In light of the lemma, it’s sufficient to define $\alpha^{\dagger}_{r}$ as follows –
(39) $p_{r_{\text{ind}}}\leq p_{r^{*}}\iff\frac{c_{r}}{1-\alpha}\leq p_{r^{*}}\iff 1-\alpha\leq\frac{c_{r}}{p_{r^{*}}}\iff\alpha\leq 1-\frac{c_{r}}{p_{r^{*}}}\equiv\alpha^{\dagger}_{r}$
Since at $\alpha=\alpha^{\dagger}_{r}$ we have $p^{\dagger}$ is well-defined per Lemma 4 (indeed, it is equal to $p_{r^{*}}$), it immediately follows that $p^{\dagger}$ is well-defined for all $\alpha>\alpha^{\dagger}_{r}$ since $\pi_{r,s}$ is strictly increasing in $\alpha$.
C.4. Relative ordering of fees
For some of the fees, the relative ordering follows immediately from the closed form expressions, for instance
(40) $\alpha_{s^{*},r^{*}}<\alpha_{r^{*}}\iff 1-\frac{c_{s}}{c_{r}}<1-\frac{c_{s}}{p_{r^{*}}}\iff c_{r}<p_{r^{*}}$
(41) $\alpha^{\dagger}_{r}<\alpha_{r^{*}}\iff 1-\frac{c_{r}}{p_{r^{*}}}<1-\frac{c_{s}}{p_{r^{*}}}\iff c_{s}<c_{r}$
As motivated above, some of the fees’ ordering depends on the seller optimal feasibility cost $c_{s^{*}}$. Indeed, we have that
(42) $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}\iff 1-\frac{c_{r}}{p_{r^{*}}}\leq 1-\frac{c_{s}}{c_{r}}\iff c_{s}\leq\frac{c_{r}^{2}}{p_{r^{*}}}\equiv c_{s^{*}}$
(43) $\alpha^{\dagger}_{r}\leq\alpha_{s^{*},r^{*}}\iff 1-\frac{c_{r}}{p_{r^{*}}}\leq 1-\frac{c_{s}}{c_{r}}\iff c_{s}\leq\frac{c_{r}^{2}}{p_{r^{*}}}=c_{s^{*}}$
Finally, we turn our attention to $\alpha^{\dagger}_{s}$ which we define as the largest fee satisfying $\pi_{r,s}(p_{s_{\text{ind}}},\alpha^{\dagger}_{s})=\pi_{r,r}(p_{r^{*}})$. Indeed, notice that we can write
$\pi_{r,s}(p_{s_{\text{ind}}}(\alpha))=\alpha p_{s_{\text{ind}}}q(p_{s_{\text{ind}}})=\left(1-\frac{c_{s}}{p_{s_{\text{ind}}}}\right)p_{s_{\text{ind}}}q(p_{s_{\text{ind}}})=(p_{s_{\text{ind}}}-c_{s})q(p_{s_{\text{ind}}})$
where we now view $p_{s_{\text{ind}}}(\alpha)=\frac{c_{s}}{1-\alpha}$ as a strictly increasing function of $\alpha$, that attains values
$p_{s_{\text{ind}}}(\alpha_{s^{*},r^{*}})=\frac{c_{s}}{1-(1-\frac{c_{s}}{c_{r}})}=c_{r};\qquad p_{s_{\text{ind}}}(1-c_{s})=1$
By continuity every price in the interval $c_{r},1]$ is attained by $p_{s_{\text{ind}}}(\alpha)$ as $\alpha$ varies from $\alpha_{s^{*},r^{*}}$ to $1-c_{s}$. With this in mind, notice that for all $p\in[c_{r},1]$ we have
$\pi_{r,r}(p)=(p-c_{r})q(p)<(p-c_{s})q(p)=\pi_{r,s}(p_{s_{\text{ind}}}(\alpha)=p)$
In particular, this implies $\pi_{r,s}(p_{s_{\text{ind}}}(\alpha))$ is a continuous function with values ranging from 0 to $\pi_{r,s}(p_{s_{\text{ind}}}(\alpha)=p_{r^{*}})>\pi_{r,r}(p_{r^{*}})$, which implies $\alpha^{\dagger}_{s}\in[0,1-c_{s}]$ is well-defined.
We can further lower bound $\alpha^{\dagger}_{s}$ by recalling from Corollary 3 that $\pi_{r,s}$ is unimodal. Thus, for any $\hat{\alpha}$ such that $\pi_{r,s}(p_{s_{\text{ind}}}(\hat{\alpha}))>\pi_{r,r}(p_{r^{*}})$, we can conclude $\alpha^{\dagger}_{s}>\hat{\alpha}$. Since by construction we have $p_{s_{\text{ind}}}(\alpha_{r^{*}})=p_{r^{*}}$, this therefore implies $\alpha^{\dagger}_{s}>\alpha_{r^{*}}$. Per (40) and (41), this further implies $\alpha^{\dagger}_{s}>\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}$, and indeed if $c_{s}\leq c_{s^{*}}$ we also have $\alpha^{\dagger}_{s}>\alpha_{s^{*}}$. By continuity of $\pi_{r,s}$, the following implication is immediate –
Corollary 0.
Suppose $\hat{\alpha}$ satisfies $\pi_{r,s}(p_{s_{\text{ind}}}(\hat{\alpha}))>\pi_{r,r}(p_{r^{*}})$. Then, for $\alpha\in[\hat{\alpha},1]$, we have $\alpha\leq\alpha^{\dagger}_{s}\iff p_{s_{\text{ind}}}\leq p^{\dagger}$.
When deriving Nash equilibria, we will find ourselves in scenarios like Corollary 5 where we have an appropriate lower bound on $\alpha$, thus for all intents and purposes we will be able to think of the relation $\alpha\leq\alpha^{\dagger}_{s}\iff p_{s_{\text{ind}}}\leq p^{\dagger}$ as true.